# The Reasoning Lingua Franca: A Double-Edged Sword for Multilingual AI

### v2

@@ Abstract
Large Reasoning Models (LRMs) achieve strong performance on mathematical, scientific, and other question-answering tasks, but their multilingual reasoning abilities remain underexplored. When presented with non-English questions, LRMs often default to reasoning in English, raising concerns about interpretability and the handling of linguistic and cultural nuances. We systematically compare an LRM’s reasoning in English versus the language of the question. Our evaluation spans two tasks: MGSM and GPQA Diamond. Beyond measuring answer accuracy, we also analyze cognitive attributes in the reasoning traces. We find that English reasoning traces exhibit a substantially higher presence of these cognitive behaviors, and that reasoning in English generally yields higher final-answer accuracy, with the performance gap increasing as tasks become more complex. However, this English-centric strategy is susceptible to a key failure mode - getting “Lost in Translation," where translation steps lead to errors that would have been avoided by question’s language reasoning.
@@ 5 Results
Figure 1 presents the final answer accuracy of LRMs on the MGSM task when reasoning in the language of the question. We find that performance is highest for English, remains strong for high-resource languages, and declines progressively for lower-resource languages (Appendix B.3.1).
Figure 2(a) compares reasoning in English with reasoning in the question’s language for the MGSM task. As we move from high-resource to low-resource languages (Appendix B.3.1), the accuracy gap between English and native-language reasoning widens, indicating that reasoning in English generally yields higher final-answer accuracy.
We next examine whether this pattern holds for the GPQA diamond task which requires expert-level domain knowledge. In Figure 3 we observe that the contrast between performance in English and non-English languages is much higher than for the MGSM task during reasoning in the question’s language. In addition, in Figure 2(b), we could see the contrast between reasoning in English and question’s language for multilingual gpqa questions is much larger than the same for MGSM (c.f. Figure 2). Notably, Danish exhibits a smaller final-answer accuracy gap between reasoning in the question’s original language and reasoning in English (c.f. Figure 2(b)). Danish performs relatively well on GPQA-Diamond compared with other non-English languages (c.f. Figure 3). This likely reflects Danish’s relatively high resource status among the languages analyzed (c.f. Appendix B.3.1). These results highlight that leveraging English is particularly important for tapping an LLM’s domain knowledge and reasoning ability on expert-level questions. Additional results are shown in appendix C.
@@ 6 Analyzing cognitive behaviors in the reasoning trace
We use gpt-4o-mini as a judge to identify and count these attributes, employing evaluation prompts (c.f. Appendix D) adapted from prior studies (Gandhi et al., 2025). For the MGSM task, we compute the average frequency of each attribute for the Qwen QwQ-32B model across languages, and report means and standard deviations over four trials.
@@ 7 Making a case for reasoning in the Question’s language
To quantify “Lost in Translation,” we measure the fraction of incorrect answers from English reasoning caused by translation mistakes; errors that would not arise if reasoning occurred directly in the question’s language. These cases are identified using gpt-4o-mini as evaluator. The evaluation setup and prompts are described in Appendix E.1. We average the proportion across four trials with the Qwen QwQ-32B model and report the standard deviation. This averaged proportion is referred to as the LiT (Lost in Translation) fraction. Appendix E.2 details the human-LM agreement study for the Lost in Translation (LiT) task.
@@ 11 Acknowledgments
We thank AI4Bharat lab members Deepon Halder, Irina Agastin, and Atharva Gawane for their assistance with annotation on the Lost in Translation task. We would also like to thank EkStep Foundation and Nilekani Philanthropies for their generous grant towards research at AI4Bharat.
@@ Appendix A Related Work - additional details
LRMs amplifies this tendency by favoring explicit reasoning in English. In this paper we present a careful, empirical comparison of reasoning performed in English versus reasoning in the question’s original language, and we highlight a seldom-discussed shortcoming of English-centric reasoning; namely, that translation-related errors can undermine correctness even when English reasoning appears stronger overall.
@@ B.1 Dataset
The MGSM dataset (Shi et al., 2022) involves 250 mathematical question-answer pairs at graduate-level for 12 languages. We extend this dataset to Hindi and Malayalam using Sarvam - Translate (SarvamAI, 2025) and to Danish using (Comanici et al., 2025). The quality of these translations are manually verified to ensure high quality.
GPQA Diamond: The GPQA Diamond dataset (Rein et al., 2024) contains 198 challenging science questions that require advanced subject expertise. It is significantly harder than the MGSM dataset. We use GPQA Diamond translated to five Indic languages; Hindi, Gujarati, Bengali, Malayalam, and Telugu (Anonymous, 2025) in this study. Additionaly, we translate GPQA Diamond dataset into Danish using (Comanici et al., 2025).
@@ C.2 Standard deviation of accuracy for MGSM
Appendix D Analyzing Cognitive Behaviors: Additional details
@@
Appendix E Lost in Translation
E.1 Lost in Translation - Quantitative Analysis
E.2 Lost in Translation: Human-LM agreement
To validate the LLM-as-a-judge evaluator, we conducted a human annotation study for "Lost in Translation" (LiT) task. We randomly sampled 50 LiT evaluation prompts from MGSM for each of three low-resource languages - Hindi, Bengali and Malayalam, yielding 150 examples in total. Each example was independently solved and labeled by native-speaker annotators with strong grade-school mathematics skills. All annotators were computer science undergraduate students and each annotated 50 questions in their respective native language. Annotators were instructed to classify a reasoning trace as “Lost in Translation” or not based on the same instructions provided to the LLM judge. We compared the LLM judgments to the human labels and computed percent agreement, precision, recall, and Cohen’s K which are detailed in Table 5. These results indicate that the LLM evaluator produces reasonably reliable annotations for the LiT task even for low resource languages like Malayalam.

### v3

@@ Abstract
Large Reasoning Models (LRMs) achieve strong performance on mathematical, scientific, and other question-answering tasks, but their multilingual reasoning abilities remain underexplored. When presented with non-English questions, LRMs often default to reasoning in English, raising concerns about interpretability and the handling of linguistic and cultural nuances. We systematically compare an LRM’s reasoning in English versus the language of the question. Our evaluation spans two tasks: MGSM and GPQA Diamond. Beyond measuring answer accuracy, we also analyze cognitive attributes in the reasoning traces. We find that English reasoning traces exhibit a substantially higher presence of these cognitive behaviors, and that reasoning in English generally yields higher final-answer accuracy, with the performance gap increasing as tasks become more complex. However, this English-centric strategy is susceptible to a key failure mode - getting “Lost in Translation," where translation steps lead to errors that would have been avoided by question’s language reasoning. Code is available at https://github.com/AI4Bharat/multilingual_reasoning_analysis.
@@ 5 Results
Figure 1 presents the final answer accuracy of LRMs on the MGSM task when reasoning in the language of the question. We find that performance is highest for English, remains strong for high-resource languages, and declines progressively for lower-resource languages.
Figure 2(a) compares reasoning in English with reasoning in the question’s language for the MGSM task. As we move from high-resource to low-resource languages, the accuracy gap between English and native-language reasoning widens, indicating that reasoning in English generally yields higher final-answer accuracy.
We next examine whether this pattern holds for the GPQA diamond task which requires expert-level domain knowledge. In Figure 3 we observe that the contrast between performance in English and non-English languages is much higher than for the MGSM task during reasoning in the question’s language. In addition, in Figure 2(b), we could see the contrast between reasoning in English and question’s language for multilingual gpqa questions is much larger than the same for MGSM (c.f. Figure 2). Notably, Danish exhibits a smaller final-answer accuracy gap between reasoning in the question’s original language and reasoning in English (c.f. Figure 2(b)). Danish performs relatively well on GPQA-Diamond compared with other non-English languages (c.f. Figure 3). This likely reflects Danish’s relatively high resource status among the languages analyzed (c.f. Section. 4.1). These results highlight that leveraging English is particularly important for tapping an LLM’s domain knowledge and reasoning ability on expert-level questions. Additional results are shown in appendix C.
@@ 6 Analyzing cognitive behaviors in the reasoning trace
We use gpt-4o-mini as a judge to identify and count these attributes, employing evaluation prompts (c.f. Appendix D) adapted from prior studies (Gandhi et al., 2025). For the MGSM task, we compute the average frequency of each attribute for the Qwen QwQ-32B model across languages, and report means and standard deviations over four trials. Appendix D.1 details the human-LLM agreement study for this task.
@@ 7 Making a case for reasoning in the Question’s language
To quantify “Lost in Translation,” we measure the fraction of incorrect answers from English reasoning caused by translation mistakes; errors that would not arise if reasoning occurred directly in the question’s language. These cases are identified using gpt-4o-mini as evaluator. The evaluation setup and prompts are described in Appendix E.1. We average the proportion across four trials with the Qwen QwQ-32B model and report the standard deviation. This averaged proportion is referred to as the LiT (Lost in Translation) fraction. Appendix E.2 details the human-LLM agreement study for the Lost in Translation (LiT) task.
@@ 11 Acknowledgments
We thank AI4Bharat lab members Deepon Halder, Irina Agastin, and Atharva Gawane for their assistance with annotation on the cognitive behavior analysis and Lost in Translation task. We would also like to thank EkStep Foundation and Nilekani Philanthropies for their generous grant towards research at AI4Bharat.
@@ Appendix A Related Work - additional details
LRMs amplify this tendency by favoring explicit reasoning in English. In this paper we present a careful, empirical comparison of reasoning performed in English versus reasoning in the question’s original language, and we highlight a seldom-discussed shortcoming of English-centric reasoning; namely, that translation-related errors can undermine correctness even when English reasoning appears stronger overall.
@@ B.1 Dataset
The MGSM dataset (Shi et al., 2022) involves 250 mathematical question-answer pairs at graduate-level for 12 languages. We extend this dataset to Hindi and Malayalam using Sarvam - Translate (SarvamAI, 2025) and to Danish using Comanici et al. (2025). The quality of these translations are manually verified to ensure high quality.
GPQA Diamond: The GPQA Diamond dataset (Rein et al., 2024) contains 198 challenging science questions that require advanced subject expertise. It is significantly harder than the MGSM dataset. In this study, we use an in-house translation of GPQA Diamond into five Indic languages: Hindi, Gujarati, Bengali, Malayalam, and Telugu. Additionally, we translate GPQA Diamond dataset into Danish using Comanici et al. (2025).
@@ C.2 Standard deviation of accuracy for MGSM
Appendix D Analyzing Cognitive Behaviors: Additional details
D.1 Human - LLM agreement study
To validate the LLM-as-a-judge evaluator, we conducted a human annotation study for the cognitive behavior analysis task. We randomly sampled 30 evaluation prompts for each of the two low-resource languages - Hindi and Malayalam, yielding 60 examples in total. Each example was independently solved and labeled by native-speaker annotators with strong grade-school mathematics skills. All annotators were computer science undergraduate students and each annotated 30 questions in their respective native language. Annotators were instructed to count and report the occurrence of each of the four cognitive behaviors in the reasoning trace using the same instructions given to the LLM evaluator. We compared the LLM judgments to the human labels and computed percent agreement and Cohen’s Kappa which are detailed in Table 4. The results indicate that the LLM evaluator delivers reasonably reliable assessments of cognitive behaviors, even for low-resource languages like Malayalam.
Appendix E Lost in Translation
E.1 Lost in Translation - Quantitative Analysis
E.2 Lost in Translation: Human-LLM agreement study
To validate the LLM-as-a-judge evaluator, we conducted a human annotation study for "Lost in Translation" (LiT) task. We randomly sampled 50 LiT evaluation prompts from MGSM for each of three low-resource languages - Hindi, Bengali and Malayalam, yielding 150 examples in total. Each example was independently solved and labeled by native-speaker annotators with strong grade-school mathematics skills. All annotators were computer science undergraduate students and each annotated 50 questions in their respective native language. Annotators were instructed to classify a reasoning trace as “Lost in Translation” or not based on the same instructions provided to the LLM judge. We compared the LLM judgments to the human labels and computed percent agreement, precision, recall, and Cohen’s Kappa which are detailed in Table 6. These results indicate that the LLM evaluator produces reasonably reliable evaluations for the LiT task even for low resource languages like Malayalam.
@@ E.2 Lost in Translation: Human-LLM agreement study
Cognitive Behavior Hindi (30 samples) Malayalam (30 samples)
% Agreement Kappa % Agreement Kappa
Subgoal 0.73 0.68 0.67 0.62
Verification 0.80 0.69 0.70 0.58
Backtracking 0.80 0.58 0.733 0.54
Backward Chaining 0.83 0.52 0.87 0.52
Table 4: Human-LLM agreement for cognitive analysis task. Percent agreement and Cohen’s Kappa are reported for Hindi and Malayalam.